# Identifying efficient routes to laminarization: an optimization approach

### v1

@@ Abstract
The inherently nonlinear and chaotic nature of turbulent flows poses a major challenge for designing effective control strategies to maintain or induce low-drag laminar states. Traditional linear methods often fail to capture the complex dynamics governing transitions between laminar and turbulent regimes. In this work, we introduce and investigate the concept of the minimal seed for relaminarization – the smallest perturbation of a turbulent state that triggers a direct transition to laminar flow without a chaotic transient. We formulate the identification of this optimal perturbation as a fully nonlinear optimization problem and develop a numerical framework based on a multi-step penalty method to compute it. Applying this framework to the nine-mode Moehlis-Faisst-Eckhardt model of a sinusoidal shear flow, we compute the minimal seeds for both transition to turbulence and relaminarization. While both of these minimal seeds lie infinitesimally close to the laminar-turbulent boundary – the so-called edge of chaos – they are generally unrelated and lie in distant and qualitatively distinct regions of state space, thereby providing different insights into the flow’s underlying structure. We find that while the optimal perturbation for triggering transition is primarily in the direction of the mode representing streamwise vortices, or rolls, the optimal perturbation for relaminarization is distributed across multiple modes in this model, without strong contributions in the roll or streak directions. While counterintuitive, this finding indicates that fully nonlinear control schemes exploiting the structure of the turbulent attractor might offer great improvements over classical methods based on linearizations and those which simply counteract the dominant features of the turbulent flow. The laminarizing trajectory obtained from the minimal seed for relaminarization also provides an efficient pathway out of the turbulent region of state space to the laminar state and can therefore inform the design and evaluation of flow control strategies for inducing laminarization.
@@ I Introduction
In this work, we approach the problem of inducing laminarization from a dynamical systems perspective by investigating the mechanisms and state-space structures underlying the relaminarization process in turbulent flows. Kerswell and coworkers [11, 12, 13] have studied the minimal seed for transition, the smallest perturbation from the laminar state that evolves toward turbulence rather than relaminarizing. Here we introduce a complementary concept, the minimal seed for relaminarization, the smallest perturbation from a turbulent attractor that leads to a laminarizing trajectory without a chaotic transient. The identification of this point can be viewed as a search for the point on the boundary between the turbulent and laminar basins of attraction that is closest to the turbulent attractor. Based on this view, we develop an optimization-based approach for identifying this minimal perturbation. The minimal seed and the laminarizing trajectory that emanates from it provide a pathway through the turbulent region of the state space to the laminar attractor, which will be useful for the development of efficient control strategies for drag reduction in turbulent flows.
@@
In many fluid flows, the laminar state is linearly stable and coexists in the high-dimensional state space with either a turbulent attractor, which sustains turbulent behavior indefinitely, or a chaotic saddle, which leads to transient turbulence before eventual laminarization [14, 15]. In these cases, an important state space structure that organizes the flow is the boundary separating laminarizing trajectories from turbulent trajectories. This boundary has been termed the edge of chaos and has been studied for numerous systems [16, 17, 18, 19, 20, 21, 22, 23]. In many of these studies, there is a particular focus on the trajectories within this boundary and the invariant solutions and state space structures which organize the dynamics within the basin boundary. Such a simple invariant solution embedded in the edge of chaos is known as an edge state. Edge states are thought to be of particular importance, as their stable and unstable manifolds have been shown to control the transition behavior between attractors in many cases [16, 18, 22]. Here, by studying the minimal seed for relaminarization and the resulting trajectory, our methods enable the identification of edge states which control the laminarization and transition processes.
@@
Many computational studies of the edge of chaos have considered low dimensional models of a fluid flow system, typically derived from a Galerkin projection of the governing equations onto a small number of modes [16, 19, 22, 21]. While these models are extreme projections of the full governing equations and do not capture the complete complexity of a turbulent fluid flow, they have proven to be useful in many cases, as they capture many aspects of the state space structure of the full flow [24, 25] and their computational efficiency allows for more involved dynamical analysis. In the present work we consider the nine-mode model of a sinusoidal shear flow introduced by Moehlis, Faisst, and Eckhardt (MFE) in Refs. [25, 26]. In this model, the laminar flow is linearly stable, so infinitesimally small perturbations always decay to laminar, and this stable laminar state coexists in the state space with turbulent solutions. For some parameter values, the turbulence is sustained, indicating the presence of a turbulent attractor. This attractor can be periodic, quasiperiodic, or chaotic depending on the domain size and Reynolds number – a bifurcation study can be found in [26, 19]. For other parameter values, the turbulence is not sustained, but only transient, with all trajectories collapsing to the laminar solution in finite time. In cases of transient turbulence, it has been shown that the distribution of turbulent lifetimes before laminarization is exponential, which is indicative of a chaotic saddle [25].
@@
The edge of chaos of the MFE model has been studied previously in Refs. [19, 21]. Kim and Moehlis [19] applied the bisection method of Skufkca et al. [16] to locate and track the edge, showing that edge trajectories for this system also tend to converge to an unstable periodic orbit, indicating that the edge of chaos coincides with the stable manifold of this periodic orbit. This periodic orbit has the smallest perturbation energy of all periodic orbits known for the system [19]. They also confirm that the transition behavior for this system is controlled by a fully nonlinear mechanism, as initial conditions which transition at small perturbation magnitudes are unrelated to the direction of maximum transient growth computed from the linearization. Joglekar et al.[21] also investigated the edge of chaos for the MFE system through visualizations on 2D projections of the trajectory lifetimes and basins of attraction for the laminar state and the chaotic attractor. Here, we show that the optimal pathway to relaminarization from the turbulent attractor in this system also passes near to the periodic orbit on the edge, indicating that this edge state controls both the laminarization and transition behavior.
@@
The minimal seed for transition is typically computed using a gradient-based optimization approach known as direct-adjoint looping, in which gradients of the optimization objective with respect to the initial condition are computed by solving an adjoint equation backward in time [28]. This procedure for computing minimal seeds was formulated by Pringle and Kerswell [11], where it was used to compute the minimal seed for transition for pipe flow. These results were extended by Pringle, et al. [12], where it was shown that trajectories on the edge near the minimal seed for transition tend to approach a travelling wave edge state, providing evidence that the minimal seed lies on the stable manifold of an edge state. Similar minimal seed computations using this optimization approach have now been carried out for a wide range of flows, including plane Couette flow [29], Poiseuille flow [30], and boundary layer flows [31, 32]. Other recent works have applied the fully nonlinear optimization approach to analyze optimal perturbations (for maximum transient energy growth) of the turbulent mean profile in a channel flow, thus shedding light on the mechanisms and coherent structures characterizing sustained turbulence [33, 34]. For the MFE system, Joglekar, et al. [21] studied the minimal perturbation magnitude needed for transition and computed the scaling of this critical perturbation with Reynolds number. In that work, however, the minimal perturbation was determined by locating the edge using the bisection strategy initialized in a large number of random initial directions, rather than the nonlinear optimization approach.
@@
In this work, we develop the related concept of the minimal seed for relaminarization, the point closest to a reference state in the turbulent region of the state space that laminarizes. From a flow control perspective, this point is of interest, as once this point in state space is reached the flow will laminarize without further actuation. Therefore, the minimal seed for laminarization and the resulting laminarizing trajectory can be seen as providing a transition pathway out of the turbulent region of the state space to the laminar solution. While the minimal seeds for relaminarization and transition both lie infinitesimally close to the edge of chaos, they are otherwise unrelated in general, and as we will see, can lie in very different regions of the state space. The minimal seeds for transition and relaminarization, along with the laminar-turbulent edge, are depicted schematically in Fig. 1.
The remainder of this paper is structured as follows. In Section II, we introduce the optimization-based approach to computing the minimal seeds for transition and relaminarization. We state the fully nonlinear optimization problem to be solved for each of these and detail the multi-step penalty method used here. We also discuss the bisection algorithm used to precisely locate the edge and track trajectories within the edge. In Section III, we detail the 9-mode model of a sinusoidal shear flow and in Section IV we present the resulting minimal seeds for transition and relaminarization for this system. While the minimal seed for transition is a streamwise vortex mode, consistent with our understanding of the self-sustaining process underlying wall turbulence, the minimal seed for relaminarization is more complex and does not involve the streamwise vortex mode at all. The potential implications of these observations for turbulence control are discussed.
II Methods
@@
Consider the governing evolution equation, written in the general form
@@
which is solved from the initial condition $\mathbf{x}(0)=\mathbf{x}_{0}$ over a time interval $t\in(0,\tau)$. We denote the stable laminar state by $\mathbf{x}_{L}$ and the distance of the trajectory from the laminar state at time $t$ by $d_{L}(t)\equiv\|\mathbf{x}(t)-\mathbf{x}_{L}\|_{2}$. With this, the constrained nonlinear program for the minimal seed for transition for a given perturbation magnitude, $d_{0}$ is
@@
For initial conditions within the basin of attraction of the laminar state, $d_{L}(\tau)$ will approach zero for sufficiently large $\tau$, but for initial conditions outside of the laminar basin of attraction, $d_{L}(\tau)$ will approach a finite value. Therefore, we can approximate the minimal seed for transition by sweeping over $d_{0}$ and searching for the smallest magnitude where this transition from $d_{L}(\tau)$ approaches a finite value. As a practical matter, the final time, $\tau$ for the optimization should be chosen to be large enough that there is a clear transition in the trajectory behavior as the perturbation magnitude increases. However, for very large $\tau$, the objective tends to become more sensitive to changes in the initial condition, which can make the optimization problem more challenging to solve.
To determine the minimal seed for relaminarization, we solve a related optimization problem. For this, we consider a reference point, $\mathbf{x}_{T}$ in the turbulent region of the state space that does not laminarize immediately. Denoting the distance from the turbulent reference at time $t$ by $d_{T}(t)=\|\mathbf{x}(t)-\mathbf{x}_{T}\|$, the constrained nonlinear optimization problem solved to determine the minimal seed for relaminarization is
@@
Solving this program over a range of $d_{0}$ and evaluating the final distance from laminar, $d_{L}(\tau)$, for the optimized initial conditions, one can determine the minimal seed for relaminarization by searching for the smallest $d_{0}$ that leads to $d_{L}(\tau)$ approaching zero for sufficiently large $\tau$. Much like the case of the minimal seed for transition, this will be indicated by a sharp transition in the value of $d_{L}(\tau)$ as $d_{0}$ is increased beyond the critical value corresponding to the distance from the turbulent reference point to the laminar-turbulent basin boundary.
@@ II.1 Multi-step penalty method
While this approach has been shown to work well in solving the minimal seed for transition problem, the minimal seed for relaminarization problem is more challenging, as it involves optimizing perturbations from a reference point, $\mathbf{x}_{T}$ that lies in the region of the state space characterized by chaotic dynamics, and therefore high sensitivity to the initial conditions. Specifically, trajectories in this chaotic region separate exponentially fast, which in turn leads to exponential growth of the gradient of the objective function with respect to the initial conditions. This property of chaotic systems makes the objective function of minimal seed for relaminarization problem highly nonconvex, which makes it difficult to employ gradient-based optimization methods to locate the true global minimum.
In light of these challenges associated with solving optimization problems over chaotic systems, Chung and Freund [35] proposed a multi-step penalty-based optimization method for solving optimization and optimal control problems in which the objective function involves the forward time-integration of a chaotic dynamical system. The idea of the multi-step method is to divide the long-time chaotic trajectory required for the optimization into many smaller segments to prevent the exponential growth of sensitivity in time from leading to prohibitively large gradients, as illustrated in Fig. 2. These shorter segments are then pieced together into one long trajectory by adding continuity constraints to the optimization problem, requiring that the terminal state of one segment matches with the initial state of the following segment. This approach has been shown to be useful in solving optimal control problems for chaotic systems [35], and more recently, for optimizing neural ODE surrogate models for chaotic systems [36].
@@
Much of the benefit of using a multi-step method is due to the fact that this formulation allows the constraints to be relaxed while solving the optimization to temporarily suppress the nonconvexity of the problem. Applying this procedure allows the solver to avoid stagnating near local extrema of the objective function by biasing the numerical solution toward regions of the state space with lower objective value before enforcing the constraints. To achieve this effect, we solve the multi-step formulation of the optimization problems using a quadratic penalty method following the approach of Ref. [35]. In a multi-step penalty method, the equality-constrained multi-step optimization is replaced by a sequence of unconstrained optimizations where the objective is augmented with a quadratic penalty for constraint violation. In the case of the minimal seed problem, only the perturbation magnitude constraint and continuity constraints will replaced by quadratic penalties. This is because the remaining constraints associated with the initial value problem can be satisfied automatically in a gradient-based optimization scheme by using an adjoint sensitivity method to compute the gradients. That is, the optimization problems are reformulated as

### v2

@@ Abstract
The nonlinear and chaotic nature of turbulent flows poses a major challenge for designing effective control strategies to maintain or induce low-drag laminar states. Traditional linear methods often fail to capture the complex dynamics governing transitions between laminar and turbulent regimes. In this work, we introduce and investigate the concept of the minimal seed for relaminarization – the closest point to a reference state in the turbulent region of the state space that triggers a direct transition to laminar flow without a chaotic transient. We formulate the identification of this optimal perturbation as a fully nonlinear optimization problem and develop a numerical framework based on a multi-step penalty method to compute it. Applying this framework to the nine-mode Moehlis-Faisst-Eckhardt model of a sinusoidal shear flow, we compute the minimal seeds for both transition to turbulence and relaminarization. While both of these minimal seeds lie infinitesimally close to the laminar-turbulent boundary – the so-called edge of chaos – they are generally unrelated and lie in distant and qualitatively distinct regions of state space, thereby providing different insights into the flow’s underlying structure. We find that while the optimal perturbation for triggering transition is primarily in the direction of the mode representing streamwise vortices, or rolls, the optimal perturbation for relaminarization is distributed across multiple modes in this model, without strong contributions in the roll or streak directions. By analyzing the trajectories originating from the minimal seeds, we find that both the transition and laminarization behavior for this system are controlled by a common mechanism – the stable and unstable manifolds of a periodic orbit on the edge of chaos. The laminarizing trajectory obtained from the minimal seed for relaminarization also provides an efficient pathway out of the turbulent region of state space to the laminar state and can therefore inform the design and evaluation of flow control strategies aimed at inducing laminarization.
@@ I Introduction
In this work, we approach the problem of inducing laminarization from a dynamical systems perspective by investigating the mechanisms and state-space structures underlying the relaminarization process in turbulent flows. Kerswell and coworkers [11, 12, 13] have studied the minimal seed for transition, the closest point to the laminar state that evolves toward the turbulent region of the state space rather than relaminarizing. Here we introduce a complementary concept, the minimal seed for relaminarization, the closest point to a reference state in the chaotic region of the state space that leads to a laminarizing trajectory without a chaotic transient. The identification of this point can be viewed as a search for the point on the boundary between the turbulent and laminar basins of attraction that is closest to the turbulent attractor. Based on this view, we develop an optimization-based approach for identifying this minimal seed. The minimal seed and the laminarizing trajectory that emanates from it provide a pathway through the turbulent region of the state space to the laminar attractor, which will be useful for the development of efficient control strategies for laminarizing chaotic systems.
@@
In many fluid flows, the laminar state is linearly stable and coexists in the high-dimensional state space with either a chaotic attractor, which sustains turbulent behavior indefinitely, or a chaotic saddle, which leads to transient turbulence before eventual laminarization [14, 15]. In these cases, an important state space structure that organizes the flow is the boundary separating laminarizing trajectories from turbulent trajectories. This boundary has been termed the edge of chaos and has been studied for numerous systems [16, 17, 18, 19, 20, 21, 22, 23]. In many of these studies, there is a particular focus on the trajectories within this boundary and the invariant solutions and state space structures which organize the dynamics within the basin boundary. Such a simple invariant solution embedded in the edge of chaos is known as an edge state. Edge states are thought to be of particular importance, as their stable and unstable manifolds have been shown to control the transition behavior between attractors in many cases [16, 18, 22]. Here, by studying the minimal seed for relaminarization and the resulting trajectory, our methods enable the identification of edge states which control the laminarization and transition processes.
@@
Many computational studies of the edge of chaos have considered low dimensional models of a fluid flow system, typically derived from a Galerkin projection of the governing equations onto a small number of modes [16, 19, 22, 21]. While these models are extreme projections of the full governing equations and do not capture the complete complexity of a turbulent fluid flow, they have proven to be useful in many cases, as they capture many aspects of the state space structure of the full flow [24, 25] and their computational efficiency allows for more involved dynamical analysis. In the present work we consider the nine-mode model of a sinusoidal shear flow introduced by Moehlis, Faisst, and Eckhardt (MFE) in Refs. [25, 26]. Although the MFE model does not truly represent fully developed turbulence in a physical flow, it captures several essential dynamical features associated with the transition to turbulence in shear flows, including the coexistence of a laminar fixed point and chaotic dynamics associated with a turbulent-like attractor. Following common usage in the literature, we use the terms “turbulent” and “laminarization” in this broader dynamical-systems sense, while recognizing that the model itself describes a reduced-order system rather than true turbulence. In this model, the laminar flow is linearly stable, so infinitesimally small perturbations always decay to laminar, and this stable laminar state coexists in the state space with turbulent solutions. For some parameter values, the turbulence is sustained, indicating the presence of a turbulent attractor. This attractor can be periodic, quasiperiodic, or chaotic depending on the domain size and Reynolds number – a bifurcation study can be found in [26, 19]. For other parameter values, the turbulence is not sustained, but only transient, with all trajectories collapsing to the laminar solution in finite time. In cases of transient turbulence, it has been shown that the distribution of turbulent lifetimes before laminarization is exponential, which is indicative of a chaotic saddle [25].
@@
The edge of chaos of the MFE model has been studied previously in Refs. [19, 21]. Kim and Moehlis [19] applied the bisection method of Skufkca et al. [16] to locate and track the edge, showing that edge trajectories for this system also tend to converge to an unstable periodic orbit, indicating that the edge of chaos coincides with the stable manifold of this periodic orbit. This periodic orbit has the smallest perturbation energy of all periodic orbits known for the system [19]. They also confirm that the transition behavior for this system is controlled by a fully nonlinear mechanism, as initial conditions which transition at small perturbation magnitudes are unrelated to the direction of maximum transient growth computed from the linearization. Joglekar et al.[21] also investigated the edge of chaos for the MFE system through visualizations on 2D projections of the trajectory lifetimes and basins of attraction for the laminar state and the chaotic attractor. Here, we show that the optimal pathway to relaminarization from the turbulent attractor in this system also passes near to the periodic orbit on the edge, indicating that this edge state controls both the laminarization and transition behavior. Lebovitz [22] studied bifurcations of the edge in the four-dimensional model of a sinusoidal shear flow from Waleffe [24], where the edge corresponds to a subset of the stable manifold of an unstable equilibrium point. This work also introduced the distinction between a strong basin boundary, which separates trajectories that never laminarize from those that do, and a weak basin boundary, where trajectories on either side laminarize and the edge marks the boundary between qualitatively different laminarization behaviors (e.g., those which laminarize immediately versus those with chaotic transience).
@@
The minimal seed for transition is typically computed using a gradient-based optimization approach known as direct-adjoint looping, in which gradients of the optimization objective with respect to the initial condition are computed by solving an adjoint equation backward in time [28]. This procedure for computing minimal seeds was formulated by Pringle and Kerswell [11], where it was used to compute the minimal seed for transition for pipe flow. These results were extended by Pringle, et al. [12], where it was shown that trajectories on the edge near the minimal seed for transition tend to approach a travelling wave edge state, providing evidence that the minimal seed lies on the stable manifold of an edge state. Similar minimal seed computations using this optimization approach have now been carried out for a wide range of flows, including plane Couette flow [29], Poiseuille flow [30], and boundary layer flows [31, 32]. Other recent works have applied the fully nonlinear optimization approach to analyze optimal perturbations (for maximum transient energy growth) of the turbulent mean profile in a channel flow, thus shedding light on the mechanisms and coherent structures characterizing sustained turbulence [33, 34]. For the MFE system, Joglekar, et al. [21] studied the minimal perturbation magnitude needed for transition and computed the scaling of this critical perturbation with Reynolds number. In that work, however, the minimal perturbation was determined by locating the edge using the bisection strategy initialized in a large number of random initial directions, rather than the nonlinear optimization approach. The nonlinear optimal perturbations for transient growth for the MFE system have also been analyzed recently by Heide and Hemati [35] through sparse optimization techniques.
@@
In this work, we develop the related concept of the minimal seed for relaminarization, defined as the point closest to a reference state in the turbulent region of the state space that laminarizes. From a control perspective, this point is useful, as once it is reached the flow will laminarize without further actuation. The minimal seed for relaminarization and the resulting laminarizing trajectory therefore provide a transition pathway out of the turbulent region of the state space. We demonstrate these ideas on the nine-mode MFE system, computing both the minimal seed for transition and the minimal seed for relaminarization, and tracking the edge from each to identify mechanism governing these transitions.
Although both minimal seeds lie infinitesimally close to the edge of chaos, they are otherwise unrelated and can lie in very different regions of the state space. In particular, the minimal seeds for relaminarization are located near the chaotic attractor on a portion of the laminar-turbulent boundary that is not visited by the edge trajectory originating from the minimal seed for transition. Relatedly, Budanur, et al. [36] also identified portions of the edge far from the laminar state in pipe flow – termed the “upper edge of chaos” there – through bisection and showed that this part of the edge can also be tracked to a common edge state for that system. We also show that the structures of the minimal seed for transition and relaminarization also differ significantly. For the MFE system, the optimal perturbation of the laminar state for transition for this system is predominantly in the direction of the streamwise vortex mode, consistent with our understanding of the self-sustaining process underlying wall turbulence. In contrast, the minimal seed for relaminarization is more complex and contains no significant component in the direction of the streamwise vortex mode. We also demonstrate how the minimal seed for relaminarization can be used as a guide for relaminarizing the flow via control. Although our analysis is carried out on the reduced-order MFE system, the formulation is general and relies on optimization-based approaches that have been applied to more complex systems. The framework developed here is intended to clarify the essential mechanisms in a setting where they can be analyzed transparently, while laying the groundwork for future extensions to more complex flow configurations. A schematic of the minimal seeds for transition and relaminarization, along with the laminar–turbulent boundary, is shown in Fig. 1.
The remainder of this paper is structured as follows. In Section II, we introduce the optimization-based approach to computing the minimal seeds for transition and relaminarization. We state the fully nonlinear optimization problem to be solved for each of these and detail the multi-step penalty method used here. We also discuss the bisection algorithm used to precisely locate the edge and track trajectories within the edge. In Section III, we detail the 9-mode model of a sinusoidal shear flow and in Section IV we present the resulting minimal seeds for transition and relaminarization for this system. The potential implications of these observations for turbulence control are discussed.
II Problem Formulation and Methods
@@
Consider the governing evolution equations, written in the general form
@@
which are solved from the initial condition $\mathbf{x}(0)=\mathbf{x}_{0}$ over a time interval $t\in(0,\tau)$. We denote the stable laminar state by $\mathbf{x}_{L}$ and the distance of the trajectory from the laminar state at time $t$ by $d_{L}(t)\equiv\|\mathbf{x}(t)-\mathbf{x}_{L}\|_{2}$. In searching for the minimal seed for transition, we seek the smallest perturbation of the laminar state that does not immediately relaminarize, but instead transitions to the chaotic region of the state space. That is, we search for the initial condition $x_{0}$ of the smallest distance from the laminar state $d_{L}(0)=\|x_{0}-x_{L}\|_{2}$ that transitions to the chaotic region before returning to the laminar state. Therefore, the search for the minimal seed for transition involves solving the following constrained optimization problem over a range of perturbation magnitudes, $d_{0}$.
@@
For initial conditions within the basin of attraction of the laminar state, $d_{L}(\tau)$ will approach zero for sufficiently large $\tau$, but for initial conditions outside of the laminar basin of attraction, $d_{L}(\tau)$ will approach a finite value. Therefore, we can approximate the minimal seed for transition by sweeping over values of $d_{0}$ and searching for the smallest magnitude where $d_{L}(\tau)$ approaches a finite value rather than zero. As a practical matter, the final time, $\tau$ for the optimization should be chosen to be large enough that there is a clear transition in the trajectory behavior as the perturbation magnitude increases. However, for very large $\tau$, the objective tends to become more sensitive to changes in the initial condition, which can make the optimization problem more challenging to solve.
To determine the minimal seed for relaminarization, we solve a related optimization problem. We consider a reference point, $\mathbf{x}_{T}$ in the turbulent region of the state space that does not laminarize immediately and denote the distance from this turbulent reference at time $t$ by $d_{T}(t)=\|\mathbf{x}(t)-\mathbf{x}_{T}\|$. To compute the minimal seed for relaminarization for a given reference point, we search for the smallest perturbation of the turbulent reference point which laminarizes without a chaotic transient. That is, we seek the initial condition, $x_{0}$, of smallest distance from the turbulent reference point, $d_{T}(0)=\|x_{0}-x_{T}\|_{2}$ such that $d_{L}(\tau)\to 0$ for sufficiently large $\tau$. Therefore, the search for the minimal seed for relaminarization involves solving the following constrained nonlinear optimization problem over a range of perturbation magnitudes, $d_{0}$.
@@
Solving this program over a range of $d_{0}$ and evaluating the final distance from laminar, $d_{L}(\tau)$, for the optimized initial conditions, one can determine the minimal seed for relaminarization by searching for the smallest $d_{0}$ that leads to $d_{L}(\tau)$ approaching zero for sufficiently large $\tau$. Much like the case of the minimal seed for transition, this will be indicated by a sharp transition in the value of $d_{L}(\tau)$ as $d_{0}$ is increased beyond the critical value corresponding to the distance from the turbulent reference point to the laminar-turbulent basin boundary. The optimal initial condition, $x_{0}$, at the critical perturbation magnitude is the minimal seed for relaminarization. In general, the minimal seed for relaminarization will depend on the chosen reference point that is being perturbed, as the nearest point on the laminar-turbulent boundary will depend on the point on the attractor that is being considered. That is, the minimal seed for relaminarization is a local property of the attractor and varies depending on the region of the attractor under consideration. For this reference point, we typically choose a point that is representative of the turbulent dynamics. Good candidates for a turbulent reference point could be fixed points, points on a periodic orbit, or points on attractor in the turbulent region. This dependence on the choice of reference point is discussed in more detail in Sec. IV.
@@ II.1 Multi-step penalty method
While this approach has been shown to work well in solving the minimal seed for transition problem, the minimal seed for relaminarization problem is more challenging, as it involves optimizing initial conditions near a reference point, $\mathbf{x}_{T}$ that lies in the region of the state space characterized by chaotic dynamics, and therefore high sensitivity to the initial conditions. Specifically, trajectories in this chaotic region separate exponentially in time, which in turn leads to exponential growth of the gradient of the objective function with respect to the initial conditions. This property of chaotic systems makes the objective function of minimal seed for relaminarization problem highly nonconvex, which makes it difficult to employ gradient-based optimization methods to locate the true global minimum.
In light of these challenges associated with solving optimization problems over chaotic systems, Chung and Freund [37] proposed a multi-step penalty-based optimization method for solving optimization and optimal control problems in which the objective function involves the forward time-integration of a chaotic dynamical system. The idea of the multi-step method is to divide the long chaotic trajectory required for the optimization into many smaller segments to prevent the exponential growth of sensitivity in time from leading to prohibitively large gradients, as illustrated in Fig. 2. These shorter segments are then pieced together into one long trajectory by adding continuity constraints to the optimization problem, requiring that the terminal state of one segment matches with the initial state of the following segment. This approach has been shown to be useful in solving optimal control problems for chaotic systems [37], and more recently, for optimizing neural ODE surrogate models for chaotic systems [38].
@@
Much of the benefit of using a multi-step method is due to the fact that this formulation allows the constraints to be relaxed while solving the optimization to temporarily suppress the nonconvexity of the problem. Applying this procedure allows the solver to avoid stagnating near local extrema of the objective function by biasing the numerical solution toward regions of the state space with lower objective value before enforcing the constraints. To achieve this effect, we solve the multi-step formulation of the optimization problems using a quadratic penalty method following the approach of Ref. [37]. In a multi-step penalty method, the equality-constrained multi-step optimization is replaced by a sequence of unconstrained optimizations where the objective is augmented with a quadratic penalty for constraint violation. In the case of the minimal seed problem, only the perturbation magnitude constraint and continuity constraints will replaced by quadratic penalties. This is because the remaining constraints associated with the initial value problem can be satisfied automatically in a gradient-based optimization scheme by using backpropagation or an adjoint sensitivity method to compute the gradients. That is, the optimization problems are reformulated as